# Accelerating Data Chunking in Deduplication Systems using Vector Instructions

### v1

@@ 5. Implementation
We accelerate AE (ae, ), MAXP (maxp, ), and RAM (ram, ) using VectorCDC with 3000 lines of C++ code. We implemented SSE-128, AVX-256, AVX-512, NEON-128, and VSX-128 versions of all algorithms. We also implemented Extreme Byte Searches for minima and maxima, as well as Range Scan functionalities with the GT, GEQ, LT, LEQ, and EQ comparators on all five vector instruction sets. We have made our code publicly available with DedupBench [^3] https://github.com/UWASL/dedup-bench (dedupbench, ).
Note that while ARM processors support VCMP and VMAX operations, they lack native support for VMASK instructions, which are used during range scans to generate a single mask containing the comparison results. This is a common issue encountered by ARM developers trying to port x86 code (armPortingVector, ). We chose an efficient alternative implementation (armPortingVector, ) to work around the lack of native VMASK support. However, this alternative implementation uses multiple slow NEON-128 instructions, such as vshrn and vreinterpretq, as opposed to a single x86 mm_movemask instruction. As shown in §6.4, this causes accelerated algorithms to achieve lower speedups on ARM CPUs compared to Intel and AMD.
@@ 6. Evaluation
Intel Emerald Rapids - -
Intel Skylake - -
AMD EPYC Rome - - -
ARM v8 Atlas - - - -
IBM Power 8 - - - -
@@
Testbed. We run all our experiments using machines from the Cloudlab (cloudlab_paper, ) platform. We pick five machines with diverse vector instruction set support; Table 1 shows the vector instruction sets supported by each machine. The details of each machine are as follows:
@@
  • CRC: Native (unaccelerated) version of the CRC-32 chunking algorithm from SS-CDC (sscdc, ).
  • FCDC: Native version of FastCDC (fastcdc, ).
  • Gear: Native version of the Gear-hash based chunking algorithm (gear_hash, ).
  • RC: Rabin’s chunking algorithm from LBFS (lbfs, ).
  • SS-CRC: AVX-512 version of CRC accelerated using SS-CDC (sscdc, ).
  • SS-Gear: AVX-512 version of Gear accelerated using SS-CDC (sscdc, ).
  • TTTD: Two-Threshold Two-Divisor algorithm (tttd, ).
@@
  • AE: Native version of the Asymmetric Extremum algorithm (ae, ). We evaluate both AE-Max and AE-Min.
  • MAXP: Native version of the MAXP algorithm (maxp, ).
  • RAM: The native Rapid Asymmetric Maximum (ram, ) algorithm.
@@
Debian (debian_org, ) VM Images obtained from
the VMware Marketplace (vmware, )
@@
DEV 230 GB 100 Nightly backups of a Rust (github_rust, ) build server 83.17% 98.05%
FLOW 8 GB 630341 Source files for 25 versions of TensorFlow (tensorflow, ) 90.69% 91.98%
KUBE 1.5 GB 117344 Source files for 5 versions of Kubernetes (kubernetes, ) 64.52% 69.42%
LNX 65 GB 160 Linux kernel distributions (kernel_linux, ) in TAR format (gnu_tar, ) 19.87% 45.62%
@@
OpenStreetMap (openstreetmap, ) backups of Canada
extracted using GeoFabrik (geofabrik, )
@@
47 consecutive days in TAR (gnu_tar, ) format
@@
RDS 122 GB 100 Redis (redis, ) snapshots between redis-benchmark runs 33.54% 92.94%
TPCC 106 GB 25 25 snapshots of a MySQL (mysql, ) VM running TPC-C (TPCCOverview, ) 37.39% 86.64%
@@
Snapshots of the largest Wikipedia article (wikipediaListFilms, ) across
@@
Datasets. We use 10 diverse datasets to evaluate VectorCDC; Table 2 shows their details. The datasets represent diverse workloads such as VM backups, database and map backups, web snapshots, and source code. Some datasets, such as FLOW and WIKI, are similar to those used by previous studies (predict_dedup_alberta_2022, ). We have publicly released the DEB dataset  [^4] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (deb_dataset, ).
We note that the selected datasets have diverse characteristics. They have varying sizes, ranging from $1$ GB for WIKI to $981$ GB for MAPS. They have different file counts; datasets such as MAPS and NEWS consist of a few large files, while others, such as FLOW and KUBE, consist of a large number of small files. We include files with varying formats, such as TAR (gnu_tar, ), VMDK / OVA (ovf_format, ), OSM (openstreetmapFileFormats, ), text files, and binary files across these datasets for comprehensive coverage.
Finally, Table 2 shows the space savings achieved by using fixed-size chunking (XC) and the median of those achieved by CDC algorithms (Median CDC) on these datasets with 8KB chunks. By comparing XC against Median CDC, we note that the datasets possess varying degrees of byte-shifting. For instance, XC achieves a space savings of 37.39% on the TPCC dataset while CDC algorithms achieve a median of 86.68%. On the other hand, XC achieves 90.69% on FLOW while CDC algorithms achieve a median of 91.98%. Two of our datasets, MAPS and WIKI, possess a large degree of byte-shifting, causing an extreme difference in space savings between XC and the median achieved by CDC algorithms.
@@
Metrics. We evaluate the space savings, chunk size distribution, and chunking throughput achieved by each alternative on all the described datasets. In addition, we examine VectorCDC’s performance across different vector instruction sets.
6.1. Space Savings and Chunk Size Distributions
Figures 8(a) - 8(j) show the space savings achieved by all alternatives with 8KB chunks across datasets. We omit the results for other chunk sizes as the trends were similar.
@@ 6.1.1. Hash-based vs Hashless
Hash-based algorithms exhibit space savings values close to each other across datasets. The best among the hashless algorithms (AE-Max, AE-Min, MAXP, and RAM) achieves slightly lower space savings than the best hash-based algorithm on some datasets, such as DEB and NEWS (Figures 8(a) and 8(f)). On the other hand, the best hashless algorithm outperforms all hash-based algorithms on other datasets, such as LNX and RDS (Figures 8(e) and 8(g)). Overall, the best hashless algorithms achieve space savings values within 6% of the best hash-based ones across all datasets and chunk sizes.
The only exception to this is MAPS (Figure 8(j)), where hashless algorithms perform worse than hash-based ones. AE-Max, the best performing hashless algorithm, achieves 11% less space savings than CRC, its best hash-based counterpart. This suggests that there are some specific datasets in which hash-based algorithms are the optimal choice. However, in most cases, hashless algorithms remain competitive with their hash-based counterparts for data deduplication.
@@ 6.1.2. Hashless algorithm comparison
The best-performing hashless algorithm depends on the dataset’s characteristics and the average chunk size. For instance, RAM achieves the highest space savings on DEB (Figure 8(a)) while MAXP does so on TPCC (Figure 8(h)). This shows that accelerating all hashless algorithms is important, as the optimal algorithm depends on the dataset’s characteristics.
(a) AE-Max / VAE-Max
(b) AE-Min / VAE-Min
(c) MAXP / VMAXP
(d) RAM / VRAM
Figure 9. Chunk size distribution CDFs of hashless algorithms and their AVX-512 accelerated versions on TPCC with an 8KB average chunk size
AE-Max either performs equivalently to or outperforms AE-Min on all datasets, except DEB and TPCC (Figures 8(a) and 8(h)). In addition, note that AE-Min is adversely affected by the byte-shifting pattern in MAPS, causing it to achieve only 8.89% in space savings while other CDC algorithms achieve 58%-78%.
@@
6.1.3. Vector-acceleration Impact
Note that SS-CRC, SS-Gear, VAE, VMAXP, and VRAM achieve the same space savings as their native counterparts, i.e., vector-acceleration does not impact the space savings achieved by CDC algorithms. This aligns with the results previously observed for SS-CRC and SS-GEAR (sscdc, ). We omit these results from Figure 8 for clarity.
6.1.4. Chunk Size Distribution
To ensure the correctness of our vector-accelerated implementations, we compared their chunk size distributions to those of their native counterparts and verified that they were equal. Figure 9 shows the chunk size distributions exhibited by AE-Max, AE-Min, MAXP, and RAM compared against their AVX-512 versions accelerated with VectorCDC. Note that each figure is a cumulative frequency (CDF) (cumulativefreq, ) plot. We use a target average chunk size of 8KB and the TPCC dataset for this experiment. The results for other datasets and chunk sizes were similar and have been omitted for clarity.
Figure 9 shows that the AVX-512 versions of each algorithm exhibit the same chunk size distribution as their unaccelerated counterparts. For instance, Figure 9(d) shows that RAM and VRAM have identical chunk size distributions. Thus, VectorCDC does not affect the chunks generated by the hashless algorithms it accelerates.
@@ 6.2. Chunking Throughput
(c) Hash-based algorithms with SS-CDC (sscdc, )
@@
Figure 10. Chunking Throughput with AVX-512 instructions and 8KB chunks. Note that the legend entries are in the same order as the plot bars from Figures 10(a) and 10(b).
@@ 6.2.1. Throughput Comparison
Figures 10(a) and 10(b) show that hashless algorithms accelerated with VectorCDC achieve the highest throughputs among all CDC algorithms. VRAM, the fastest accelerated hashless algorithm, achieves $8.35\times$ and $15.3\times$ higher throughput than SS-GEAR and FastCDC, the fastest accelerated and unaccelerated hash-based algorithms, respectively.
Among unaccelerated hash-based algorithms, FastCDC (fastcdc, ), Gear (gear_hash, ), and CRC (sscdc, ) are the fastest, achieving 1.95 GB/s, 1.05 GB/s, and 0.48 GB/s, respectively. We accelerated each of these using SS-CDC (sscdc, ); SS-GEAR achieves 3.57 GB/s and SS-CRC achieves 1.17 GB/s. We discuss accelerating FastCDC in §6.2.2.
The unaccelerated hashless algorithms AE-Max (ae, ), AE-Min (ae, ), MAXP (maxp, ), and RAM (ram, ) come in at $1.5$ GB/s–$1.7$ GB/s, faster than most of their hash-based counterparts. Their VectorCDC-accelerated versions, VAE-Max, VAE-Min, VMAXP, and VRAM, achieve $6.5$ GB/s–$29.9$ GB/s. VRAM, the fastest accelerated hashless algorithm, is $8.35\times$ and $26.2\times$ faster than SS-GEAR and SS-CRC respectively. While the throughputs of all algorithms slightly vary across datasets, the trends remain the same.
6.2.2. Accelerating FastCDC.
We did not observe any speedup when accelerating FastCDC (fastcdc, ) with SS-CDC (sscdc, ). One of the main throughput optimizations used by FastCDC is sub-minimum skipping (§2.1). However, as noted in §3.2, decoupling the rolling-hash phase from the boundary identification phase eliminates the throughput benefits of minimum chunk size skipping, nullifying any speedup provided by vector-acceleration. VAE-Max, VAE-Min, VMAXP and VRAM achieve $6\times, 4.4$$46\times$, $4.76\times$, and $15.3\times$ higher throughput than FastCDC respectively.
@@
Accelerating hash-based algorithms (Figure 10(c)) using SS-CDC achieves a speedup of $2.45-3.32\times$. For instance, on DEB, SS-CRC exhibits a throughput of $1.2$ GB/s over CRC at $0.46$ GB/s, i.e., a $2.45\times$ speedup. Similarly, SS-GEAR achieves $3.57$ GB/s over GEAR at $1.07$ GB/s, i.e., a $3.32\times$ speedup.
On the other hand, the hashless algorithms VAE-Max, VAE-Min, VMAXP, and VRAM achieve speedups of $5.1\times$, $4.43\times$, $5.36\times$, and $17.69\times$ over their respective native counterparts. Thus, vector instructions can be leveraged far more efficiently for hashless algorithms, proving that hashless algorithms are better candidates for vector-acceleration than their hash-based counterparts.
@@
Figure 11 shows the time taken by the chunking and hashing phases in the deduplication pipeline on DEB with an 8KB average chunk size. We omit the results for other datasets as they were similar. We used xxHash3, the fastest hashing algorithm (§3.1) for fingerprinting and ran this experiment on the Intel Emerald Rapids machine. We use AVX-512 versions of hashless CDC algorithms, accelerated with VectorCDC. Note that the times in Figure 3 and Figure 11 differ due to the differing sizes and dataset characteristics between the random dataset used in §3 and DEB.
Figure 11. Time taken for Data Chunking vs Fingerprinting on DEB with an 8 KB average chunk size, xxHash3 for fingerprinting, and AVX-512 instructions for acceleration.
For unaccelerated CDC algorithms, Figure 11 shows similar results to those obtained in Figure 3 with random data. Data chunking takes significantly longer than fingerprinting with these algorithms. On the other hand, VAE-Min, VAE-Max, VMAXP, and VRAM show data chunking times similar to or lower than fingerprinting. For instance, with VRAM, data chunking takes $1.29$ seconds while fingerprinting takes $2.27$ seconds.
This shows that VectorCDC effectively alleviates the data chunking bottleneck in the deduplication pipeline.

### v2

@@ 5. Implementation
We accelerate AE (Zhang et al., 2015), MAXP (Bjørner et al., 2010), and RAM (Widodo et al., 2017) using VectorCDC with 3000 lines of C++ code. We implemented SSE-128, AVX-256, AVX-512, NEON-128, and VSX-128 versions of all algorithms. We also implemented Extreme Byte Searches for minima and maxima, as well as Range Scan functionalities with the GT, GEQ, LT, LEQ, and EQ comparators on all five vector instruction sets. We have made our code publicly available with DedupBench [^3] https://github.com/UWASL/dedup-bench (Liu et al., 2023).
Note that while ARM processors support VCMP and VMAX operations, they lack native support for VMASK instructions, which are used during range scans to generate a single mask containing the comparison results. This is a common issue encountered by ARM developers trying to port x86 code (Kutenin, 2022). We chose an efficient alternative implementation (Kutenin, 2022) to work around the lack of native VMASK support. However, this alternative implementation uses multiple slow NEON-128 instructions, such as vshrn and vreinterpretq, as opposed to a single x86 mm_movemask instruction. As shown in §6.4, this causes accelerated algorithms to achieve lower speedups on ARM CPUs compared to Intel and AMD.
@@ 6. Evaluation
Intel Emerald Rapids
Intel Skylake
AMD EPYC Rome
ARM v8 Atlas
IBM Power 8
@@
Testbed. We run all our experiments using machines from the Cloudlab (Duplyakin et al., 2019) platform. We pick five machines with diverse vector instruction set support; Table 1 shows the vector instruction sets supported by each machine. The details of each machine are as follows:
@@
  • CRC: Native (unaccelerated) version of the CRC-32 chunking algorithm from SS-CDC (Ni et al., 2019).
  • FCDC: Native version of FastCDC (Xia et al., 2016b).
  • Gear: Native version of the Gear-hash based chunking algorithm (Xia et al., 2014).
  • RC: Rabin’s chunking algorithm from LBFS (Muthitacharoen et al., 2001).
  • SS-CRC: AVX-512 version of CRC accelerated using SS-CDC (Ni et al., 2019).
  • SS-Gear: AVX-512 version of Gear accelerated using SS-CDC (Ni et al., 2019).
  • TTTD: Two-Threshold Two-Divisor algorithm (Eshghi and Tang, 2005).
@@
  • AE: Native version of the Asymmetric Extremum algorithm (Zhang et al., 2015). We evaluate both AE-Max and AE-Min.
  • MAXP: Native version of the MAXP algorithm (Bjørner et al., 2010).
  • RAM: The native Rapid Asymmetric Maximum (Widodo et al., 2017) algorithm.
@@
Debian (Debian, 2025) VM Images obtained from
the VMware Marketplace (VMWare, 2023)
@@
DEV 230 GB 100 Nightly backups of a Rust (Rust, 2023) build server 83.17% 98.05%
FLOW 8 GB 630341 C++ source code for 25 versions of TensorFlow (Pang et al., 2020) 90.69% 91.98%
KUBE 1.5 GB 117344 Go source code for 5 versions of Kubernetes (Luksa, 2017) 64.52% 69.42%
LNX 65 GB 160 Linux kernel distributions (Linux, 2023) in TAR format (GNU, 2023) 19.87% 45.62%
@@
OpenStreetMap (Haklay and Weber, 2008) backups of Canada
extracted using GeoFabrik (GeoFabrik, 2025)
@@
47 consecutive days in TAR (GNU, 2023) format
@@
RDS 122 GB 100 Redis (Redis, 2023) snapshots between redis-benchmark runs 33.54% 92.94%
TPCC 106 GB 25 25 snapshots of a MySQL (MySQL, 2023) VM running TPC-C (Council, 2023) 37.39% 86.64%
@@
Snapshots of the largest Wikipedia article (Wikipedia, 2022) across
@@
Datasets. We use 10 diverse datasets to evaluate VectorCDC; Table 2 shows their details. The datasets represent diverse workloads such as VM backups, database and map backups, web snapshots, and source code. Some datasets, such as FLOW and WIKI, are similar to those used by previous studies (Randall and Lu, 2022). We have publicly released the DEB dataset  [^4] https://www.kaggle.com/datasets/sreeharshau/vm-deb-fast25 (Udayashankar et al., 2025b).
We note that the selected datasets have diverse characteristics. They have varying sizes, ranging from $1$ GB for WIKI to $981$ GB for MAPS. They have different file counts; datasets such as MAPS and NEWS consist of a few large files, while others, such as FLOW and KUBE, consist of a large number of small files. We include files with varying formats, such as OSM (OpenStreetMap, 2025), RDB (Rediger, 2015), TAR (GNU, 2023), VMDK / OVA (DMTF, 2009), text files, and binary files across these datasets for comprehensive coverage.
Finally, Table 2 shows the space savings achieved by using fixed-size chunking (XC) and the median of those achieved by CDC algorithms (Median CDC) on these datasets with 8KB chunks. By comparing XC against Median CDC, we note that the datasets possess varying degrees of byte-shifting. The difference in space savings between XC and Median CDC in FLOW and KUBE is small (less than 6% ), indicating a smaller number of byte-shifts. DEV has a moderate amount of byte-shifting, as shown by the $15\%$ difference between XC and Median CDC. Finally, CDC algorithms achieve a median of more than $2\times$ higher space savings than XC on DEB, LNX, MAPS, NEWS, RDS, TPCC, and WIKI, indicating that these data sets have a large degree of byte-shifting.
(a) AE-Max / VAE-Max
(b) AE-Min / VAE-Min
(c) MAXP / VMAXP
(d) RAM / VRAM
Figure 8. Chunk size CDFs of hashless algorithms and their AVX-512 accelerated versions on TPCC with an 8KB average chunk size
Metrics. We evaluate the space savings, chunk size distribution, and chunking throughput achieved by each alternative on all the described datasets.
6.1. Space Savings and Chunk Size Distributions
Figures 9(a) - 9(j) show the space savings achieved by all alternatives with 8KB chunks across datasets. We omit the results for other chunk sizes as the trends were similar.
6.1.1. Vector-acceleration Impact
Vector-acceleration does not impact the space savings achieved by CDC algorithms. Consequently, for clarity, we omit the space savings results for vector-accelerated algorithms from Figure 9. This aligns with the results previously observed for SS-CRC and SS-GEAR (Ni et al., 2019).
AVX-512 acceleration does not impact the chunks generated by hashless algorithms. We compared the generated chunks of vector-accelerated algorithms with their native counterparts and verified that they were identical. We present only the chunk size distribution comparison in this paper due to space constraints.
Figure 8 shows the chunk size distributions exhibited by AE-Max, AE-Min, MAXP, and RAM compared against their AVX-512 versions accelerated with VectorCDC. Note that each figure is a cumulative frequency (CDF) (Burr, 1942) plot. We use a target average chunk size of 8KB and the TPCC dataset for this experiment. The results for other datasets and chunk sizes were similar and have been omitted for clarity.
@@ 6.1.2. Hash-based vs Hashless
Hashless algorithms are generally competitive with hash-based ones in space savings. The best among the hashless algorithms achieves slightly lower space savings than the best hash-based algorithm on some datasets, such as DEB and NEWS (Figures 9(a) and 9(f)). On the other hand, the best hashless algorithm outperforms all hash-based algorithms on other datasets, such as LNX and RDS (Figures 9(e) and 9(g)). Overall, the best hashless algorithms achieve space savings values within 11% of the best hash-based ones across all datasets and chunk sizes.
@@ 6.1.3. Hashless algorithm comparison
The performance of the hashless algorithms depends on the dataset’s characteristics and the average chunk size. For instance, RAM achieves the highest space savings on DEB (Figure 9(a)) while MAXP does so on TPCC (Figure 9(h)). This shows that accelerating all hashless algorithms is important, as the performance of each algorithm depends on the dataset’s characteristics.
Notably, AE-Min is adversely affected by the byte-shifting pattern in MAPS, causing it to achieve only 8.89% in space savings while other CDC algorithms achieve 58%-78%.
@@
6.1.4. Differences among datasets.
Hashless algorithms perform equivalent to or better than their counterparts on virtual machine and database backups, such as DEV, RDS, and TPCC. Source-code datasets demonstrate mixed results, with hash-based algorithms slightly edging out hashless ones on KUBE, equivalence on FLOW, and hashless algorithms being better on LNX. File formats largely do not influence space savings.
@@ 6.2. Chunking Throughput
(c) Hash-based algorithms with SS-CDC (Ni et al., 2019)
@@
Figure 10. Chunking Throughput with AVX-512 instructions and 8KB chunks. Note the different scales in Figures 10(c) and 10(d), and that the legend entries are in the same order as the plot bars from Figures 10(a) and 10(b).
@@ 6.2.1. Throughput Comparison
Figures 10(a) and 10(b) show that hashless algorithms accelerated with VectorCDC achieve $4\times$ to $15\times$ higher throughput than all accelerated CDC algorithms. VRAM, the fastest accelerated hashless algorithm, achieves $8.35\times$ and $15.3\times$ higher throughput than SS-GEAR and FastCDC, the fastest accelerated and unaccelerated hash-based algorithms, respectively. Additionally, VRAM achieves $207.2\times$ higher throughput than RC, a popular but slow hash-based CDC algorithm.
Among unaccelerated hash-based algorithms, Gear (Xia et al., 2014), CRC (Ni et al., 2019), and FastCDC (Xia et al., 2016b) are the fastest. We accelerated each of these using SS-CDC (Ni et al., 2019); SS-GEAR achieves 3 $\times$ higher throughput compared to its unaccelerated version, and SS-CRC achieves 2 $\times$ higher throughput that unaccelerated CRC. We did not observe any speedup when accelerating FastCDC (Xia et al., 2016b) with SS-CDC (Ni et al., 2019). One of the main throughput optimizations used by FastCDC is sub-minimum skipping (§2.1). However, as noted in §3.2, decoupling the rolling-hash phase from the boundary identification phase eliminates the throughput benefits of minimum chunk size skipping, nullifying any speedup provided by vector-acceleration.
@@
Accelerating hash-based algorithms (Figure 10(c)) using SS-CDC achieves a speedup of $2.45-3.32\times$. On the other hand, the hashless algorithms VAE-Max, VAE-Min, VMAXP, and VRAM achieve speedups of $5.1\times$, $4.43\times$, $5.36\times$, and $17.69\times$ over their respective native counterparts, achieving throughputs in the range of $6.5$ GB/s–$29.9$ GB/s. Thus, vector instructions can be leveraged far more efficiently for hashless algorithms, proving that hashless algorithms are better candidates for vector-acceleration than their hash-based counterparts.
(a) xxHash3 - 128-bit digest
(b) SHA256 - 256-bit digest
Figure 11. Time taken for Data Chunking vs Fingerprinting on DEB with an 8 KB average chunk size, and AVX-512 instructions for acceleration.
@@
Figure 11 shows the time taken by the chunking and hashing phases in the deduplication pipeline on DEB with an 8KB average chunk size. We omit the results for other datasets as they were similar. We used two fingerprinting algorithms; xxHash3, the fastest but generates a 128-bit digest, and SHA-256, slower but offers higher collision resistance with a 256-bit digest3.1). We ran this experiment on the Intel Emerald Rapids machine. We use AVX-512 versions of hashless CDC algorithms, accelerated with VectorCDC.
Figure 11 shows that with xxHash3 (Figure 11(a)), data chunking takes significantly longer than fingerprinting with unaccelerated algorithms. On the other hand, VAE-Min, VAE-Max, VMAXP, and VRAM show data chunking times similar to or lower than fingerprinting. For instance, with VRAM, data chunking takes $1.29$ seconds while fingerprinting takes $2.27$ seconds.
With SHA-256 (Figure 11(b)), we observe that fingerprinting takes as long as data chunking with unaccelerated CDC algorithms. On the other hand, VAE-Max, VAE-Min, VMAXP, and VRAM take significantly lower time to run.
These results show that VectorCDC effectively alleviates the data chunking bottleneck in the deduplication pipeline.